# Adaptive Beam Alignment using Noisy Twenty Questions Estimation with Trained Questioner

### v1

@@ V Conclusion
There are several other avenues for future directions. Firstly, in this paper, we proposed two methods for adaptive BA, aiming to simultaneously achieve feasibility, interpretability and low latency. However, there remains a gap between their performance and the theoretical optimum. One can explore other non-NN or NN methods to narrow this gap, potentially even reaching the theoretical limits. Secondly, we focused on the LoS setting in this paper, i.e., there is only one AoA to be estimated. But non-LoS setting is more common in practice, involving multiple AoAs to be estimated. In that case, the optimal beamforming vector should exhibit a multi-cluster structure [21]. It is thus of interest to extend the active learning framework to the multi-target case, derive the corresponding theoretical results and apply it to the non-LoS setting. Thirdly, we only considered the sortPM strategy proposed in [6, Eq. (13)] in this paper, while introducing alternative strategies, such as hiePM and dyaPM, may yield new insights and better performance. It would also be interesting to investigate the cases where sortPM is replaced with hiePM and dyaPM in our proposed algorithms. Furthermore, we followed the same strategy as in [6, Eq. (13)] to generate the queries. It may lead to a mismatch between the actual distribution of the generated codebook and the capacity-achieving input distribution, which can degrade the algorithms’ performance. It is worthwhile to find out how to adjust the strategy to eliminate the mismatch, or how to quantify the performance loss if the mismatch is unavoidable. Finally, we focused on the stationary target in this paper. However, a target in motion represents a more realistic setting in communication systems. Our results for BA of a stationary target can be generalized to the moving target case via ideas in [23, 35, 33, 25].

### v2

@@ V Conclusion
There are several other avenues for future directions. Firstly, in this paper, we proposed two methods for adaptive BA, aiming to simultaneously achieve feasibility, interpretability and low latency. However, there remains a gap between their performance and the theoretical optimum. One can explore other non-NN or NN methods to narrow this gap, potentially even reaching the theoretical limits. Secondly, we focused on the LoS setting in this paper, i.e., there is only one AoA to be estimated. But non-LoS setting is more common in practice, involving multiple AoAs to be estimated. In that case, the optimal beamforming vector should exhibit a multi-cluster structure [21]. It is thus of interest to extend the active learning framework to the multi-target case, derive the corresponding theoretical results and apply it to the non-LoS setting. Thirdly, we only considered the sortPM strategy proposed in [6, Eq. (13)] in this paper, while introducing alternative strategies, such as hiePM and dyaPM, may yield new insights and better performance. It would also be interesting to investigate the cases where sortPM is replaced with hiePM and dyaPM in our proposed algorithms. Furthermore, we followed the same strategy as in [6, Eq. (13)] to generate the queries. It may lead to a mismatch between the actual distribution of the generated codebook and the capacity-achieving input distribution, which can degrade the algorithms’ performance. It is worthwhile to find out how to adjust the strategy to eliminate the mismatch, or how to quantify the performance loss if the mismatch is unavoidable. Finally, we focused on the stationary target case in this paper. However, a target in motion represents a more realistic setting in communication systems. Our results of the adaptive BA for a stationary target can be generalized to the moving target case via ideas in [23, 35, 33, 25].